# Local Stability and Region of Attraction Analysis for Neural Network Feedback Systems under Positivity Constraints

### v1

@@ I Introduction
Neural networks (NNs) are increasingly being used in control systems due to their expressive power and ability to learn complex nonlinear mappings[1]. However, certifying the stability of NN-in-the-loop systems remains a major challenge, particularly when such systems are deployed in safety-critical applications. Traditional Lyapunov-based analysis is often infeasible for high-dimensional or opaque NN controllers, motivating the need for scalable, structure-exploiting methods for stability verification.
@@
Existing techniques that attempt to verify stability for NN feedback—such as those using integral quadratic constraints (IQCs)[2, 3], sum-of-squares optimization [4], passivity-based approaches [5], circle criterion [6] or global Lipschitz bounds [7]—either scale poorly with system size or result in overly conservative guarantees. Recent advances such as RecurJac[8], CROWN [9], Bnb [10] and IBP [11] provide tools for bounding NN outputs via layer-wise analysis, but these have not been explicitly integrated into closed loop system frameworks.
@@
Positive systems offer a favorable structure for a streamlined and scalable stability analysis. These systems have been widely explored in the literature of linear systems [12]. However, their benefits in nonlinear systems have not been extensively studied. The positive Lur’e framework provides a well-established setting for studying absolute stability of complex nonlinear systems. For example, the positive Aizerman conjecture gives sufficient conditions for global exponential stability in this context when the feedback nonlinearity lies within a sector [13]. However, there is currently a lack of techniques that can leverage the structural properties of positive systems to enable scalable and less conservative verification of NN feedback stability.
@@
This paper addresses this gap by proposing two tractable and scalable methods for local stability analysis and estimating the ROA of positive Lur’e systems with NN feedback, both grounded in a localized version of the positive Aizerman conjecture. The first method uses a Lyapunov-based approach with linear matrix inequalities (LMIs), applicable to general sector-bounded nonlinearities. The second introduces a novel, tight local sector bound for FFNNs, which to the best of our knowledge, is the first such formulation in the literature. This bound is propagated layer by layer through the network and naturally fits within the Aizerman framework, enabling a scalable and certifiable ROA analysis for NN feedback systems. Compared to existing bounds such as RecurJac, CROWN, and IBP, this novel local sector bound—built upon our previously proposed global sector bound [14]—features a nonaffine linear shape, and is specifically developed for scalable local stability analysis of NN feedback loops.
The key contributions of this paper are:
  • A localized variant of the Positive Aizerman Conjecture tailored for local stability analysis and further calculating the corresponding ROA;
  • A Lyapunov-based method for local stability analysis and ROA estimation in positive Lur’e systems with general sector-bounded feedback;
  • A novel local sector bound formulation for FFNNs, and further integration of it in a local stability analysis framework for NN feedback loops.
@@
These contributions are significant since they provide a formal guarantee that despite the nonlinear nature of the NN, there is a specific area in the state space where the system is stabilizable. This result highlights the practical utility of NN controllers in real-world applications, where local stability is often sufficient. Through rigorous analysis and simulation, this paper provides both theoretical insights and practical guidelines for implementing NN controllers in feedback systems. This opens up new possibilities for using machine learning techniques in classical control theory, where traditional linear control methods may fall short [15].
@@ I-A Notations
We denote the set of real numbers, real vectors of dimension $n$, and real matrices of dimension $n\times m$ by $\mathbb{R},\mathbb{R}^{n}$, and $\mathbb{R}^{n\times m}$. The orders >,<,>,<$\geq$> , < , ≥, and $\leq$ are interpreted elementwise when applied to vectors and matrices. In addition, the operators $\succ$ and $\prec$ denote positive and negative definite matrices. Subsequently, a doubly positive matrix $A$ is both $A\succ 0$ and $A>0$. The symbol $\mathbf{0}$ denotes a zero vector of appropriate dimension. We also define $\mathbb{R}_{+}^{n\times m}:=\{A\in\mathbb{R}^{n\times m}|A\geq 0\}$. The sets $\mathbb{R}_{+}^{n}$ and $\mathbb{R}_{+}$ are defined similarly. A Metzler matrix is characterized by having nonnegative off-diagonal elements.
@@ II Proposed Theoretical Framework
First, we lay the theoretical foundations for presenting our results, encompassing positive Lur’e systems and the positive Aizerman conjecture. The following definition and lemmas along with their proof can be found in [16].
@@ II-A Positive Lur’e Systems
There exists a unique, locally absolutely continuous function $\chi:\mathbb{R}_{+}\to\mathbb{R}^{n}$ that satisfies (2) almost everywhere for any $x_{0}\in\mathbb{R}^{n}$. Moreover, $\Phi(0,t)=0,\forall t\geq 0$, ensuring that $x_{*}=0$ is an equilibrium point of (2). The function $\Phi(z,t)$ is locally Lipschitz in $z$ and measurable in $t$, with certain mild boundedness assumptions [17].
Furthermore, to enable stability analysis, it is necessary to characterize the sector bounds of the nonlinear component. For a multi-input multi-output (MIMO) positive system, these bounds are best described using component-wise inequalities. The function $\Phi$ is said to be sector bounded in $[\Sigma_{1},\Sigma_{2}]$ if:
@@ II-B Positive Aizerman Conjecture
Consider a Lur’e system in (2), where $B,C\geq 0,$ and $\Phi$ is sector bounded in the interval $[\Sigma_{1}$, $\Sigma_{2}]$ in the sense of (3). The system is globally exponentially stable if $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz.
@@ II-C Problem Formulation
Consider the Lur’e system as defined in (2), with $B,C\geq 0,$ and the equilibrium point $x_{*}$ satisfying $\mathbf{0}=Ax_{*}+Bu_{*},\;y_{*}=Cx_{*},\;u_{*}=\Phi(y_{*})$. Let $\chi(t;x_{0})$ denote the solution to the closed-loop system at time $t$, initialized at $x_{0}$. Our goal is to analyze the stability of the equilibrium point $x_{*}$, and to identify a maximal underapproximation of the set of initial conditions from which the system trajectories converge to this point. This set is commonly referred to as the ROA in nonlinear systems literature and is defined as: :=$\mathcal{R}:=\{x_{0}\in\mathbb{R}^{n}:\lim_{t\to\infty}\chi(t;x_{0})=x_{*}\}.$
@@ III Local Stability via Positive Aizerman Conjecture
Given an interval $[\Sigma_{1},\Sigma_{2}]$ with $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{m\times p}$ and $\Sigma_{1}\leq\Sigma_{2}$, a multivariate nonlinear function $\Phi(y,\cdot):\mathbb{R}^{p}\times\mathbb{R}\to\mathbb{R}^{m}$ is said to be $\Gamma-$sector bounded within that interval if, for a compact and connected set $\Gamma\subseteq\mathbb{R}^{p}$, the following condition holds:
@@
Consider the Lur’e system given in (2), where $B,C\geq 0$ and the nonlinearity $\Phi$ is $\Gamma-$sector bounded within the interval $[\Sigma_{1},\Sigma_{2}]$ with $y_{*}\in\Gamma$. Suppose that $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz; then all the output trajectories that remain within the set $\Gamma$ are exponentially stable.
@@
Consider a multi-input single-output (MISO) Lur’e system satisfying theorem 2. Let $\overline{y}$ denote the maximal element of the set $\Gamma$ in which the nonlinearity is sector bounded. Then, an underapproximation of the ROA is given by the set of initial conditions $x_{0}$ satisfying:
@@
The Lyapunov function presented in the proof of Lemma 3 with equations (6) and (7), alongside the result of [13, Lemma 2], provides a proof for our Theorem 2.
@@ IV Lyapunov Based Method
Theorem 3.
Consider the Lur’e system formulated in (2), with $B,C\geq 0$. Let the nonlinearity $\Phi$ be $\Gamma$-sector bounded over the interval $[\Sigma_{1},\Sigma_{2}]$. If the matrix $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz, then there exists a Lyapunov function that certifies the local stability of the nonlinear system. Specifically, a verified Lyapunov function is given by $V(x)=x^{\top}Px$, where $P\in\mathbb{R}^{n\times n}$ is a doubly positive matrix satisfying the linear matrix inequality:
@@
The first part of the theorem follows by invoking the converse Lyapunov theorem, which guarantees the existence of a Lyapunov function for locally asymptotically stable systems. In particular, by Theorem 2, the Lur’e system under consideration is locally asymptotically stable. Thus, there exists a continuously differentiable, positive definite function $V(x)$ such that its derivative along system trajectories satisfies $\dot{V}(x)<0$ in a neighborhood of the origin.
Now consider the quadratic Lyapunov candidate $V(x)=x^{\top}Px$, where $P$ is a doubly positive matrix satisfying the LMI in (10). Define the time derivative of $V$ along system trajectories:
@@
By the assumption that the nonlinearity $\Phi$ is $\Gamma-$sector bounded in the interval $[\Sigma_{1},\Sigma_{2}]$, we obtain the inequality:
@@
for some neighborhood $\mathcal{D}$ of the origin. Noting that $B$, $P$, and $x$ are all nonnegative elementwise and using the monotonicity of scalar multiplication under the elementwise inequality (12), we deduce:
$2x^{\top}PB\Phi(Cx,.)\leq 2x^{\top}PB\Sigma_{2}Cx.$ (13)
@@
By construction of $P$, the matrix inequality (10) implies that the right-hand side is strictly negative. Hence, $\dot{V}(x)<0$ for all $x\neq 0$ in some neighborhood of the origin, completing the proof of local asymptotic stability via the Lyapunov function $V(x)=x^{\top}Px$. ∎
@@
1:Lur’e system matrices $A,B>0,C>0$, nonlinearity $\Phi(y,.)$ $\Gamma$-sector bounded in $[\Sigma_{1},\Sigma_{2}]$
@@
This method is applicable to general nonlinearities, including neural networks. It significantly reduces the search space for ROA from arbitrary sets to ellipsoidal regions aligned with Lyapunov levels. By focusing on invariant sublevel sets rather than the global sector region $\Gamma$, the approach mitigates the conservatism typically associated with worst-case sector analysis and provides a practically verifiable estimate of the region of attraction.
@@ V Local Sector Bounds for FFNN
In this section, we revisit the Aizerman conjecture as a foundational principle for the local stability analysis of nonlinear feedback systems. Given that the conjecture fundamentally relies on the existence of sector bounds for the system nonlinearity, we begin by introducing a novel formulation of local sector bounds for NNs specifically tailored to the stability analysis theorems. In particular, we propose a new bounding framework for FFNNs without bias terms. To the best of our knowledge, this is the first sector based characterization of its kind within the machine learning literature.
While related bounds have been explored in the context of robustness and verification—such as the product of weight norms [7], the CROWN framework [19], RecurJac [8], and local Lipschitz estimates [20]—our proposed method offers several key advantages. Unlike the norm-based upper bounds in [7], our construction yields significantly tighter enclosures. Moreover, in contrast to CROWN [19], our formulation avoids the inclusion of affine offset terms, making it particularly compatible with the linear sector condition required in the Aizerman conjecture framework.
Furthermore, compared to Lipschitz-based relaxations [20, 8], our approach fixes the slope of the linear relaxation of the activations at the origin—rather than taking the derivative of the activation function at the boundaries of pre-activation logits—thereby reducing computational overhead while preserving fidelity near the equilibrium. This property significantly enhances scalability, allowing the technique to be applied efficiently to high-dimensional networks without the need for costly symbolic or gradient-based computations.
We consider fully connected FFNNs without biases, although the proposed approach is readily extendable to non-fully connected architectures with minor modifications. The network structure is defined as follows:
@@
$\nu^{(i)}=W^{(i)}\omega^{(i-1)},$ $i=1,\dots,q$ (16b)
$\omega^{(i)}=\phi^{(i)}(\nu^{(i)}),$ $i=1,\dots,q$ (16c)
$u=W^{(q+1)}\omega^{(q)},$ (16d)
where the input is denoted by $y\in\mathbb{R}_{+}^{n_{0}}$, and the output of the $i$-th layer (pre-activation) is given by $\nu^{(i)}\in\mathbb{R}^{n_{i}}$ for $i=1,\dots,q$. Moreover, $W^{(i)}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ is the weight matrix at layer $i$, and $\phi^{(i)}(\cdot)$ denotes the elementwise activation function applied at that layer.

### v2

@@ I Introduction
Neural networks (NNs) are increasingly being used in control systems due to their expressive power and ability to learn complex nonlinear mappings[1]. There are recent interesting usage of NN structures in Control [2, 3]. However, certifying the stability of NN-in-the-loop systems remains a major challenge, particularly when such systems are deployed in safety-critical applications. Traditional Lyapunov-based analysis is often infeasible for high-dimensional or opaque NN controllers, motivating the need for scalable, structure-exploiting methods for stability verification.
@@
Existing techniques that attempt to verify stability for NN feedback—such as those using integral quadratic constraints (IQCs)[4, 5, 6, 7], sum-of-squares optimization [8], passivity-based approaches [9], circle criterion [10] or global Lipschitz bounds [11]—either scale poorly with system size or result in overly conservative guarantees. Recent advances such as RecurJac [12], CROWN [13], and IBP [14] provide tools for bounding NN outputs, but these have not been explicitly integrated into closed loop system frameworks. There are other useful control methods such as density functions [15] or positive system methods [16] that has not yet extended into verification of NN in the loop.
@@
Positive systems offer a favorable structure for a streamlined and scalable stability analysis. These systems have been widely explored in the literature of linear systems [17, 18]. However, their benefits in nonlinear systems have not been extensively studied. The positive Lur’e framework provides a well-established setting for studying absolute stability of complex nonlinear systems. For example, the positive Aizerman conjecture gives sufficient conditions for global exponential stability in this context when the feedback nonlinearity lies within a sector [19]. However, there is currently a lack of techniques that can leverage the structural properties of positive systems to enable scalable and less conservative verification of NN feedback stability.
@@
This paper addresses this gap by proposing two tractable and scalable methods for local stability analysis and estimating the ROA of positive Lur’e systems with NN feedback, both grounded in a localized version of the positive Aizerman conjecture. The first method uses a Lyapunov-based approach with linear matrix inequalities (LMIs), applicable to general sector-bounded nonlinearities. The second introduces a novel, tight local sector bound for FFNNs, which to the best of our knowledge, is the first such formulation in the literature. This bound is propagated layer by layer through the network and naturally fits within the Aizerman framework, enabling a scalable and certifiable ROA analysis for NN feedback systems. Compared to existing bounds such as RecurJac, CROWN, and IBP, this novel local sector bound—built upon our previously proposed global sector bound [20] and its application in robust stability of NN feedback loops [21]—features a nonaffine linear shape, and is specifically developed for scalable local stability analysis of NN feedback loops. The key contributions of this paper are: (i) A localized variant of the Positive Aizerman Conjecture tailored for local stability analysis and further calculating the corresponding ROA; (ii) A Lyapunov-based method for local stability analysis and ROA estimation in positive Lur’e systems with general sector-bounded feedback; (iii) A novel local sector bound formulation for FFNNs, and further integration of it in a local stability analysis framework for NN feedback loops.
@@
These contributions are significant since they provide a formal guarantee that despite the nonlinear nature of the NN, there is a specific area in the state space where the system is stabilizable. This result highlights the practical utility of NN controllers in real-world applications, where local stability is often sufficient. Through rigorous analysis and simulation, this paper provides both theoretical insights and practical guidelines for implementing NN controllers in feedback systems. This opens up new possibilities for using machine learning techniques in classical control theory, where traditional linear control methods fall short [22, 23].
@@ I-A Notations
We denote the set of real numbers, real vectors of dimension $n$, and real matrices of dimension $n\times m$ by $\mathbb{R}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{n\times m}$. The orders $>,<,\geq$, and $\leq$ are interpreted element-wise when applied to vectors and matrices. In addition, the operators $\succ$ and $\prec$ denote positive and negative definite matrices. Subsequently, a doubly positive matrix $A$ is both $A\succ 0$ and $A>0$. The symbol $\mathbf{0}$ denotes a zero vector of appropriate dimension. We also define $\mathbb{R}_{+}^{n\times m}:=\{A\in\mathbb{R}^{n\times m}\mid A\geq 0\}$. The sets $\mathbb{R}_{+}^{n}$ and $\mathbb{R}_{+}$ are defined similarly. A Metzler matrix is characterized by having nonnegative off-diagonal elements.
@@ II Proposed Theoretical Framework
First, we lay the theoretical foundations for presenting our results, encompassing positive Lur’e systems and the positive Aizerman conjecture. throughout the paper, we briefly use positivity for internal positivity. The following definition and lemmas along with their proof can be found in [24].
@@ II-A Positive Lur’e Systems
There exists a unique, locally absolutely continuous function $\chi:\mathbb{R}_{+}\to\mathbb{R}^{n}$ that satisfies (2) almost everywhere for any $x_{0}\in\mathbb{R}^{n}$. Moreover, $\Phi(0,t)=0,\forall t\geq 0$, ensuring that $x_{*}=0$ is an equilibrium point of (2). The function $\Phi(z,t)$ is locally Lipschitz in $z$ and measurable in $t$, with certain mild boundedness assumptions [25, Thm. 54, Prop. C.3.8].
Furthermore, to enable stability analysis, it is necessary to characterize the sector bounds of the nonlinear component. For a multi-input multi-output (MIMO) positive system, these bounds are best described using component-wise inequalities. The function $\Phi$ is said to be sector bounded in $\Sigma_{1},\Sigma_{2}]$ if:
@@ II-B Positive Aizerman Conjecture
Consider a Lur’e system in (2), where $B,C\geq 0,$ and $\Phi$ is sector bounded in the interval $\Sigma_{1}$, $\Sigma_{2}]$ in the sense of (3). The system is globally exponentially stable if $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz.
@@ II-C Problem Formulation
Consider the Lur’e system as defined in (2), with $B,C\geq 0,$ and the equilibrium point $x_{*}$ satisfying $\mathbf{0}=Ax_{*}+Bu_{*}$, $y_{*}=Cx_{*}$, $u_{*}=\Phi(y_{*})$. Let $\chi(t;x_{0})$ denote the solution to the closed-loop system at time $t$, initialized at $x_{0}$. Our goal is to analyze the stability of the equilibrium point $x_{*}$, and to identify a maximal underapproximation of the set of initial conditions from which the system trajectories converge to this point. This set is commonly referred to as the ROA in nonlinear systems literature and is defined as: $\mathcal{R}:=\{x_{0}\in\mathbb{R}^{n}:\lim_{t\to\infty}\chi(t;x_{0})=x_{*}\}.$
@@ III Local Stability via Positive Aizerman Conjecture
Given an interval $\Sigma_{1},\Sigma_{2}]$ with $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{m\times p}$ and $\Sigma_{1}\leq\Sigma_{2}$, a multivariate nonlinear function $\Phi(y,\cdot):\mathbb{R}^{p}\times\mathbb{R}\to\mathbb{R}^{m}$ is said to be $\Gamma$-sector bounded within that interval if, for a compact and connected set $\Gamma\subseteq\mathbb{R}^{p}$, the following condition holds:
@@
Consider the Lur’e system given in (2), where $B,C\geq 0$ and the nonlinearity $\Phi$ is $\Gamma$-sector bounded within the interval $\Sigma_{1},\Sigma_{2}]$ with $y_{*}\in\Gamma$. Suppose that $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz; then all the output trajectories that remain within the set $\Gamma$ are exponentially stable.
@@
Consider a multi-input single-output (MISO) Lur’e system satisfying theorem 2. Let $\overline{y}$ denote the upper bound of the set $\Gamma$ in which the nonlinearity is sector bounded. Then, an underapproximation of the ROA is given by the set of initial conditions $x_{0}$ satisfying:
@@
The Lyapunov function presented in the proof of Lemma 3 with equations (6) and (7), alongside the result of [19, Lem. 2], provides a proof for our Theorem 2.
@@ IV Lyapunov Based Method
Theorem 3 (Quadratic certificate).
For the Lur’e system (2) with $B,C\geq 0$ and a nonlinearity $\Phi$ that is $\Gamma$-sector bounded on $\Sigma_{1},\Sigma_{2}]$, if $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz, then there exists a doubly positive $P\in\mathbb{R}^{n\times n}$ such that
@@
Consequently, $V(x)=x^{\top}Px$ is a Lyapunov function certifying local stability; moreover, for some $\alpha>0$, the sublevel set $\mathcal{V}_{\alpha}=\{x:V(x)\leq\alpha\}$ is compact, positively invariant, and forms an inner approximation of the ROA.
@@
Consider the quadratic Lyapunov candidate $V(x)=x^{\top}Px$, where $P$ is a doubly positive matrix satisfying the LMI in (10). Define the time derivative of $V$ along system trajectories:
@@
By the assumption that the nonlinearity $\Phi$ is $\Gamma$-sector bounded in the interval $\Sigma_{1},\Sigma_{2}]$, we obtain the inequality:
@@
for some neighborhood $\mathcal{D}$ of the origin. Noting that $B$, $P$, and $x$ are all nonnegative element-wise and using the monotonicity of scalar multiplication under the element-wise inequality (12), we deduce:
$\small 2x^{\top}PB\Phi(Cx,.)\leq 2x^{\top}PB\Sigma_{2}Cx.$ (13)
@@
By construction of $P$, the matrix inequality (10) implies that the right-hand side is strictly negative. Hence, $\dot{V}(x)<0$ for all $x\neq 0$ in some neighborhood of the origin, completing the proof of local asymptotic stability. ∎
@@
1:Lur’e system matrices $A,B>0,C>0$, nonlinearity $\Phi(y,.)$ $\Gamma$-sector bounded in $\Sigma_{1},\Sigma_{2}]$
@@
This method is applicable to general nonlinearities, including NNs. It significantly reduces the search space for ROA from arbitrary sets to ellipsoidal regions aligned with Lyapunov levels. By focusing on invariant sublevel sets rather than the global sector region $\Gamma$, the approach mitigates the conservatism typically associated with worst-case sector analysis and provides a practically verifiable estimate of the region of attraction.
@@ V Local Sector Bounds for FFNN
Given that the positive Aizerman conjecture fundamentally relies on the existence of sector bounds for the system nonlinearity, we begin by introducing a novel formulation of local sector bounds for NNs specifically tailored to the stability analysis theorems. To the best of our knowledge, this is the first sector based characterization of its kind within the machine learning literature.
Relative to other available bounds—norm products [11], CROWN [13], RecurJac [12], and local Lipschitz estimates [26]—our construction (i) is tighter than norm-based bounds, (ii) avoids CROWN’s affine offsets, aligning with the linear sector condition in the Aizerman framework, and (iii) fixes sector slopes at the origin rather than taking derivative of the activation function at boundaries of pre-activation logits, cutting computational overhead while preserving fidelity near equilibrium. The method scales to high-dimensional networks without symbolic or gradient calculations. We focus on fully connected FFNNs without biases, although these constraints can be relaxed with a straightforward extension of the method. The network structure is defined as follows:
@@
$\nu^{(i)}=W^{(i)}\omega^{(i-1)},\qquad\omega^{(i)}=\phi^{(i)}(\nu^{(i)}),\qquad i=1,\dots,q$ (16b)
$u=W^{(q+1)}\omega^{(q)},$ (16c)
where the input is denoted by $y\in\mathbb{R}_{+}^{n_{0}}$, and the output of the $i$-th layer (pre-activation) is given by $\nu^{(i)}\in\mathbb{R}^{n_{i}}$ for $i=1,\dots,q$. Moreover, $W^{(i)}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ is the weight matrix at layer $i$, and $\phi^{(i)}(\cdot)$ denotes the element-wise activation function applied at that layer.